# An extrapolated and provably convergent algorithm for nonlinear matrix decomposition with the ReLU function

### v1

@@ 5.2 EDMC
[…]
Figure 3: Euclidean distance matrix completion. Distribution of points on the left and relative error with the Euclidean distance matrix of the points in the right. The $x$-axis represents the percentage of known entries.
We observe from Figure 3 that the distribution of the points affects the recovery performance of the algorithms:
  • When the points are randomly distributed, $30\%$ of the entries needs to be observed for the eBCD-NMD to achieve an error below $10^{-7}$.
  • When the points are clustered, $50\%$ of the entries needs to be observed for the eBCD-NMD to achieve an error below $10^{-7}$.
Including the rank-1 modification in the model allows for a significantly more accurate reconstruction than solving a rank-($r+1$) 3B-ReLU-NMD. In general, A-NMD, eBCD-NMD and 3B-NMD provide the best reconstructions. However, 3B-NMD fails at least once in both examples when the percentage of observed entries is above $80\%$. This behavior might be due to the aggressive extrapolation procedure that 3B-NMD uses. We emphasize that neither the 3B-ReLU-NMD nor the Latent-ReLU-NMD formulations impose any constraints to ensure that the solution is symmetric. However, if the number of observed entries is sufficiently large, the algorithms still recover the original symmetric solution.
@@ 5.3 Low-dimensional embedding
In this subsection, we test eBCD-NMD within the threshold similarity matching (TSM) framework of Saul [33] for dimensionality reduction. Specifically, ReLU-NMD is employed to find a lower-dimensional representation of the similarity matrix of the input points ; see Section 2 for the details. Our focus is on text data, where each row of the data matrix corresponds to a document, each column to a word, and the entries $(i,j)$ of the matrix is the number of occurrences of the $j$th word in the $i$th document. We consider two different well-known and already pre-pocessed text-word data sets from [44], that are k1b which has 2340 documents and 21839 words and hitech that has 2301 documents and 10080 words. The threshold parameter $\tau$ in the TSM framework is selected according to the statistical analysis proposed in [33], that yields $\tau=0.17$ for k1b and $\tau=0.08$ for hitech.
Our objective is to evaluate how eBCD-NMD performs in comparison to the state-of-the-art algorithms for computing ReLU decomposition in this context. The quality of the embedding is evaluated using the mean angular deviation (MAD) metric as suggested in [33]. The MAD measures the difference between the angles that the algorithm tries to preserve (namely the ones below $\cos^{-1}(\tau)$). Define $\mathcal{P}(\tau)=\{(i,j)\ |\ \langle z_{i},z_{j}\rangle-\tau\lVert z_{j}\rVert\lVert z_{i}\rVert>0\}$, the MAD is given by
@@
where the $z_{i}$’s are the higher-dimensional points, and the $y_{i}$’s the lower-dimensional embeddings. Figure 4 displays the MAD and average running time per iteration for increasing values of the approximation rank. We run all the algorithms with a fixed time limit of 60 seconds.
Figure 4: MAD analysis and average iteration time for increasing values of the rank in the $x$-axes. k1b dataset on top and hitech dataset at the bottom.
Even though we are not able to identify one algorithm that outperforms the others, Figure 4 shows that solving 3B-ReLU-NMD is to be preferred in this application. Indeed, BCD-NMD, eBCD-NMD and 3B-NMD achieve the lowest MAD in both experiments. In fact, these three algorithms have the lowest average iteration time, being more than one second faster per iteration than Naive-NMD, A-NMD and EM-NMD. For this reason, BCD-NMD, eBCD-NMD and 3B-NMD perform more iterations than Naive-NMD, A-NMD and EM-NMD in 60 seconds, resulting in more accurate approximations.
@@ 5.4 Compression of sparse data
The last application is the compression of sparse data. We consider sparse greyscale images: 10000 randomly selected images from MNIST and Fashion MNIST (fMNIST), both of size $784\times 10000$, and on two well-known sparse images, Phantom ($256\times 256$) and Satellite ($256\times 256$). We also use sparse, structured matrices from various applications from [10]: Trec11 ($235\times 1138$), mycielskian ($767\times 767$), lock ($1074\times 1074$), and beaconfd ($173\times 295$).
We choose the approximation rank $r$ to ensure a desired level of compression. Specifically, when computing the ReLU factors $W$ and $H$ of a sparse matrix $X\in\mathbb{R}^{m\times n}$ with $\operatorname{nnz}(X)$ non zero entries, the number of entries of $WH$ is $r(n+m)$. Hence, the rank $r$ is set to the closest integer such that $r\leq 0.5\cdot\operatorname{nnz}(X)/(n+m)$, which means we reduced the memory requirement by $50\%$. We compare eBCD-NMD with state-of-the-art methods for ReLU decomposition, and we compute the TSVD of the same rank as baseline.
Table 2 shows the relative error $\frac{\|X-\max(0,\bar{W}\bar{H})\|_{F}}{\|X\|_{F}}$, where $\bar{W}$ and $\bar{H}$ are the factors of the rank-$r$ approximation produced by any algorithm. We run each algorithm for 120 seconds and display the average error over 10 different random initializations.
Data set BCD-NMD eBCD-NMD 3B-NMD Naive-NMD A-NMD EM-NMD TSVD
MNIST 12.4% 11.8% 11.6% 12.5% 11.7% 13.2% 25.8%
fMNIST 9.7% 9.3% 9.3% 9.7% 9.2% 9.9% 14.0%
Phantom 5.1% 4.1% 4.3% 5.4% 4.5% 5.6% 19.2%
Satellite 14.0% 13.6% 13.7% 14.4% 14.0% 14.8% 26.0%
Trec11 27.4% 26.5% 26.5% 27.2% 26.3% 31.1% 57.3%
mycielskian 0.37% 0.05% 0.06% 0.53% 0.06% 3.6% 58.0%
lock1074 0.3% 0.001% 0.001% 1.5% 0.006% 66.7% 67.2%
beaconfd 20.6% 20.4% 20.0% 20.9% 20.7% 26.8% 39.1%
Table 2: Sparse matrices compression. Comparison of the ReLU decomposition error in percentage on various data sets. The best values are highlighted in bold.
@@
We observe that ReLU models outperform the linear alternative TSVD, as all the algorithms yield approximations that are significantly more accurate than TSVD. Additionally, eBCD-NMD and 3B-NMD produce the most accurate solutions in seven out of eight examples; moreover, eBCD-NMD outperforms BCD-NMD, achieving a smaller compression error for all the data sets considered. We also observe that the structure of the matrix has an impact on the effectiveness of the compression using ReLU decomposition.
Finally, we propose two visual examples of $50\%$ compression for MNIST and Phantom in Figure 5, where we compare TSVD with the approximation $\max(0,\Theta)$ provided by the eBCD-NMD algorithm. In both instances, one can visually appreciate how the eBCD-NMD approximation is consistently more accurate than TSVD, as further confirmed by the average error reported in Table 2.
Original digits
TSVD compression
eBCD compression
Original
TSVD compression
eBCD compression
Figure 5: Comparison between $50\%$ compressed images using TSVD and eBCD on MNIST and Phantom data sets.
@@ 6 Conclusion
The ReLU decomposition is a relatively recent matrix decomposition model (Saul, 2022 [33]) that finds applications in the compression of sparse data, entry-dependent matrix completion, and manifold learning. Moreover, we showed, for the first time, how it can be used for the completion of Euclidean distance matrices when only the smallest (or largest) entries are observed. Our first main contribution is to establish the non-equivalence of f the two main models used to solve the ReLU problem, namely ReLU-NMD and Latent-ReLU-NMD (Corollary 3.5). Furthermore, we derive an explicit connection between the optimal function values of the two, demonstrating that the solution of Latent-ReLU-NMD is in the worst case twice larger than that of ReLU-NMD (Theorem 3.6). Our second main contribution is to prove convergence of the BCD-NMD scheme for solving 3B-ReLU-NMD by demonstrating that it falls into the general framework of exact BCD, where one of the subproblems is strictly quasi-convex (Theorem 4.1). Our third and most important contribution is a novel algorithm, eBCD-NMD, which uses extrapolation to accelerate eBCD-NMD. eBCD-NMD extends the well-known LMaFit algorithm from [43] for matrix completion. We proved subsequence convergence of the eBCD-NMD scheme to a stationary point under mild assumptions (Theorem 4.9). Finally, we provided extensive numerical results for 4 applications on synthetic and real data sets, showing that eBCD-NMD performs favorably compared to the to the state of the art, while having convergence guarantees which other algorithms lack.
Future work includes investigating the well-posedness of the standard ReLU-NMD in (1), designing randomized algorithms to tackle large-scale problems, and exploring more advanced ReLU decompositions such as $X\approx W_{1}H_{1}+\max(0,W_{2}H_{2})-\max(0,W_{3}H_{3})$ that could handle mixed-signed and non-sparse data.
@@ Acknowledgments
G.S. and N.G. acknowledge the support by the European Union (ERC consolidator, eLinoR, no 101085607). The work of M.P. and G.S. was partially supported by INdAM-GNCS under the project CUP E53C23001670001. The research of M.P.  was partially granted by the Italian Ministry of University and Research (MUR) through the PRIN 2022 “MOLE: Manifold constrained Optimization and LEarning”, code: 2022ZK5ME7 MUR D.D. financing decree n. 20428 of November 6th, 2024 (CUP B53C24006410006).
@@ 7 Appendix
Similarly, we want to write the second term of the summation in terms of the residual. We use (41) to get
$H(\alpha)=$ $W(\alpha)^{T}Z_{\alpha}=W(\alpha)^{T}(WH+\alpha S)=$
@@
Let us first show the continuity of the function in (28) when $\alpha\to 1^{+}$. Recall that $Z_{1}=Z$ and that $W(\alpha)$ is an orthogonal basis of $\mathcal{R}(Z_{\alpha}H^{T})$. Therefore, the orthogonal projection operator into $\mathcal{R}(Z_{\alpha}H^{T})$ can be written as $W(\alpha)W(\alpha)^{T}=(Z_{\alpha}H^{T})(Z_{\alpha}H^{T})^{\dagger}$. If rank($Z_{\alpha}H^{T}$)=rank($ZH^{T}$) in some interval $[1,\widehat{\alpha}]$, then by the continuity of the Moore-Penrose inverse [37], we have
@@
$=$ $\lVert P_{\Omega^{C}}(W(1)H(1))\rVert_{F}^{2}+\lVert P_{\Omega^{C}}(\min(0,WH))\rVert_{F}^{2}-2\langle P_{\Omega^{C}}(W(1)H(1)),P_{\Omega^{C}}(\min(0,WH))\rangle-$
@@
$-2\langle P_{\Omega^{C}}(\min(0,W(1)H(1))),P_{\Omega^{C}}(\min(0,WH))\rangle-2\langle P_{\Omega^{C}}(\max(0,W(1)H(1))),P_{\Omega^{C}}(\min(0,WH))\rangle$
$=$ $\lVert P_{\Omega^{C}}(\min(0,W(1)H(1))-(\min(0,WH))\rVert_{F}^{2}-2\langle P_{\Omega^{C}}(\max(0,W(1)H(1))),P_{\Omega^{C}}(\min(0,WH))\rangle\geq 0,$

### v2

@@ 5.2 EDMC
We generate 200 points in two different ways in a three-dimensional space: (1) uniformly at random in $0,10]^{3}$, and (2) randomly clustered around six clusters (4 with 30 points, and two with 40) where the centroid $(\bar{x},\bar{y},\bar{z})$ of the $i$th cluster in sampled uniformly at random in $10,10]^{3}$, and we generate the points in MATLAB using $(\bar{x},\bar{y},\bar{z})+\text{std}*\texttt{randn(1,3)}$, where std denotes the standard deviation that we fixed to std=3. We compute the Euclidean distance matrix $\Theta^{t}$ containing the squared distance between all couples of points and we construct our target matrix $X=\max(0,dee^{T}-\Theta^{t})$, where $d$ represents the largest value in the matrix that we want to observe. The upper bound on the rank is fixed to $r=5$ which is equal to the rank of $\Theta^{t}$. In this experiment, we assume the threshold $d$ to be known, but one might be interested in considering $d$ as an unknown of the model. We aim to recover all the entries of $\Theta^{t}$ larger than $d$. In practice, the threshold $d$ is adapted to achieve the percentage of observed entries in the $x$-axis of the pictures on the right of Figure 2. Specifically, we display the average relative error of the low-rank approximation with the Euclidean distance matrix $\Theta^{t}$ for different percentages of observed entries, over 10 random experiments. Moreover, all algorithms have been adapted to include the rank-1 modification in (5). We excluded the EM algorithm from the analysis because adapting it to the rank-1 modified model appears less straightforward than for the other algorithms. We also consider as a baseline the rank-$(r+1)$ reconstruction given by the original BCD method solving the 3B-RMD of rank $r+1$, without considering any rank-1 modification (($r+1$)-BCD). In this case, we compute a rank-($r+1$) decomposition of $X$ and evaluate the relative error with $dee^{T}-\Theta^{t}$. We stop the algorithms when a tolerance of $10^{-9}$ on the relative residual $\Gamma_{k}$ is reached or after 60 seconds.
Figure 2: Euclidean distance matrix completion: relative errors where the $x$-axis represents the percentage of known entries. The left figure is for the uniform distribution of the data points, the right figure for the clustered distribution.
The distribution of the points affects the recovery performance of the algorithms, see Figure 2: When the points are randomly distributed, $30\%$ of the entries need to be observed for the eBCD to achieve an error below $10^{-7}$. When the points are clustered, only $50\%$ of the entries need to be observed for the eBCD to achieve an error below $10^{-7}$. Including the rank-1 modification in the model allows for a significantly more accurate reconstruction than solving a rank-($r+1$) 3B-RMD. In general, A-Naive, eBCD, and e3B provide the best reconstructions. However, e3B fails at least once in both examples when the percentage of observed entries is above $80\%$. This behavior might be due to the aggressive extrapolation procedure that e3B uses. We emphasize that neither the 3B-RMD nor the Latent-RMD formulations impose any constraints to ensure that the solution is symmetric. However, if the number of observed entries is sufficiently large, the algorithms still recover the original symmetric solution.
@@ 5.3 Low-dimensional embedding
RMD can be employed to find a lower-dimensional representation of the similarity matrix of the input points; see Section 2.4. Our focus is on text data, where each row of the data matrix corresponds to a document, each column to a word, and the entries $(i,j)$ of the matrix is the number of occurrences of the $j$th word in the $i$th document. We consider two well-known and already pre-pocessed text-word data sets from [zhong2005generative]: k1b with 2340 documents and 21839 words, and hitech with 2301 documents and 10080 words. The threshold parameter $\tau$ in the TSM framework is selected according to the statistical analysis proposed in [saul2022geometrical], that is, $\tau=0.17$ for k1b and $\tau=0.08$ for hitech. The quality of the embedding is evaluated using the mean angular deviation (MAD) metric as suggested in [saul2022geometrical]:
@@
where the $z_{i}$’s are the higher-dimensional points, and the $y_{i}$’s the lower-dimensional embeddings, and $\mathcal{P}(\tau)=\{(i,j)\ |\ \langle z_{i},z_{j}\rangle-\tau\lVert z_{j}\rVert\lVert z_{i}\rVert>0\}$, Figure 3 displays the MAD and average run times per iteration for different ranks. We run all the algorithms for 60 seconds.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 3: MAD analysis and average iteration time for increasing values of the rank: The top two figures are for the k1b dataset, and the bottom two for the hitech dataset.
Even though no algorithm outperforms the others, Figure 3 shows that using the 3B-RMD is to be preferred in this application. Indeed, BCD, eBCD and e3B achieve the lowest MAD in both experiments. Moreover, these three algorithms have the lowest average iteration time, being more than one second faster per iteration than Naive, A-Naive, and EM. For this reason, BCD, eBCD, and e3B perform more iterations than Naive, A-Naive, and EM within 60 seconds, resulting in more accurate approximations.
@@ 5.4 Compression of sparse data
The last application is the compression of sparse data. We consider sparse greyscale images: 10000 randomly selected images from MNIST and Fashion MNIST (fMNIST), both of size $784\times 10000$, and on two well-known sparse images: Phantom ($256\times 256$) and Satellite ($256\times 256$). We also use sparse, structured matrices from various applications from [dai2012geometric]: Trec11 ($235\times 1138$), mycielskian ($767\times 767$), lock ($1074\times 1074$), and beaconfd ($173\times 295$). We choose the approximation rank $r$ to ensure a desired level of compression. Specifically, when computing the RMD factors $W$ and $H$ of a sparse matrix $X\in\mathbb{R}^{m\times n}$ with $\operatorname{nnz}(X)$ non zero entries, the number of entries of $WH$ is $r(n+m)$. Hence, the rank $r$ is set to the closest integer such that $r\leq 0.5\cdot\operatorname{nnz}(X)/(n+m)$, which means a compression of $50\%$. Table 2 shows the relative error $\frac{\|X-\max(0,\bar{W}\bar{H})\|_{F}}{\|X\|_{F}}$, where $\bar{W}$ and $\bar{H}$ are the computed factors. Since we are considering real-world data of different dimensions, using a single time limit that is suitable for all the dataset is not meaningful. Therefore, we assign a time limit to all the algorithms (except the TSVD) scaled by the dimensions of each data set: we set a maximum time limit of $T=300$ seconds for the largest data sets, namely MNIST and fMNIST with $m_{\max}=784$ and $n_{\max}=10000$. Then, we assign a scaled time limit $T_{\text{d}}$ for a data set of dimension $m_{\text{d}}\times n_{\text{d}}$ using $T_{\text{d}}=T\frac{m_{\text{d}}\times n_{\text{d}}}{m_{\max}\times n_{\max}}$ seconds, and display the average relative error over 10 different random initializations.
Data set BCD eBCD e3B Naive A-Naive EM TSVD
MNIST (300 s.) err 12.2% 11.6% 11.5% 12.3% 11.6% 12.9% 25.8%
$r=70$ iter 1215 2159 1075 1088 899 413 /
fMNIST (300 s.) err 9.6% 9.1% 9.2% 9.5% 9.1% 9.7% 14.0%
$r=181$ iter 703 1498 637 914 777 513 /
Phantom (2.5 s.) err 9.0% 6.4% 7.1 9.6% 7.8% 9.8% 19.2%
$r=26$ iter 540 2898 508 374 340 278 /
Satellite (2.5 s.) err 16.1% 14.9% 15.0% 17.0% 15.6% 18.2% 26.0%
$r=12$ iter 1052 4385 983 374 352 252 /
Trec11 (10.2 s.) err 31.3% 28.9% 28.8% 30.6% 28.8% 35.7% 57.3%
$r=13$ iter 526 902 398 691 461 228 /
mycielskian (22.5 s.) err 3.6% 0.6% 0.9% 8.0% 1.3% 9.0% 58.0%
$r=14$ iter 516 1021 400 232 213 106 /
lock1074 (44.2 s.) err 6.3% 0.1% 0.2% 15.6% 1.9% 29.2% 67.3%
$r=12$ iter 576 1158 432 238 214 96 /
beaconfd (2.0 s.) err 22.8 % 22.0% 21.7% 23.3% 21.7% 37.3% 39.12%
$r=3$ iter 998 1514 680 525 419 206 /
Table 2: Compression of sparse matrices: we compare the average final relative error and the average iteration number, within the given time limit. The best values are highlighted in bold.
@@
We observe that RMD outperform the TSVD, as all the algorithms yield approximations that are significantly more accurate than the TSVD. Additionally, eBCD produces the most accurate solutions in five out of eight examples; moreover, eBCD outperforms BCD, achieving a smaller compression error for all the data sets considered. We also observe that the structure of the matrix has a crucial impact on the effectiveness of the compression using RMD. For example, for lock1074, the improvement of RMD is impressive: the relative error of eBCD is 0.1% while that of the TSVD is 67.3%. On the other hand, the improvement for fMNIST is less pronounced: from 14% relative error for the TSVD to 9.6% for RMD.
Figure 4 visually illustrates the difference between TSVD and RMD for $50\%$ compression. One can visually appreciate that the eBCD approximation is consistently more accurate than the TSVD, as hinted by the errors reported in Table 2.
Refer to caption Refer to caption Refer to caption
Original images TSVD eBCD
Figure 4: Visual illustration: TSVD (relative error = 25.8%) against RMD with eBCD (relative error = 11.6%) on the MNIST set with $50\%$ compression ($r=70$).
@@ 6 Conclusion
The RMD is a relatively recent matrix decomposition model (Saul, 2022 [saul2022geometrical]) that finds applications in the compression of sparse data, entry-dependent matrix completion, and manifold learning. Moreover, we showed, for the first time, how it can be used for the completion of Euclidean distance matrices when only the smallest (or largest) entries are observed. Our first main contribution is to establish the non-equivalence of the two main models used to to compute RMDs, namely LS-RMD and Latent-RMD (Corollary 3.5). Furthermore, we derive an explicit connection between the optimal function values of the two, demonstrating that the solution of Latent-RMD is, in the worst case, twice larger than that of LS-RMD (Theorem 3.6). Our second main contribution is to prove the convergence of the BCD scheme for solving 3B-RMD by demonstrating that it falls into the general framework of exact BCD, where one of the subproblems is strictly quasi-convex (Theorem 4.1). Our third and most important contribution is a novel algorithm, eBCD, which uses extrapolation to accelerate BCD. Our new algorithm extends the well-known LMaFit algorithm from [wen2012solving] for matrix completion. We proved the subsequence convergence of the eBCD-NMD scheme to a stationary point under mild assumptions (Theorem 4.9). Finally, we provided extensive numerical results for 4 applications on synthetic and real data sets, showing that eBCD performs favorably compared to the to the state of the art, while having convergence guarantees which other algorithms lack.
Future work includes investigating the well-posedness of the LS-RMD in (1), designing randomized algorithms to tackle large-scale problems, and exploring more advanced RMDs such as $X\approx W_{1}H_{1}+\max(0,W_{2}H_{2})-\max(0,W_{3}H_{3})$ that could handle mixed-signed and non-sparse data.
@@ Acknowledgments
We are grateful to the anonymous reviewers who carefully read the manuscript, their feedback helped us improve our paper.
@@ 7 Appendix
Similarly, we write the second term of the summation in terms of the residual. Using (39),
$H(\alpha)=$ $W(\alpha)^{T}Z_{\alpha}=W(\alpha)^{T}(WH+\alpha S)$
@@
Let us first show the continuity of the function in (26) when $\alpha\to 1^{+}$. Recall that $Z_{1}=Z$ and that $W(\alpha)$ is an orthogonal basis of $\mathcal{R}(Z_{\alpha}H^{T})$. Therefore, the orthogonal projection operator into $\mathcal{R}(Z_{\alpha}H^{T})$ can be written as $W(\alpha)W(\alpha)^{T}=(Z_{\alpha}H^{T})(Z_{\alpha}H^{T})^{\dagger}$. If rank($Z_{\alpha}H^{T}$)=rank($ZH^{T}$) in some interval $1,\widehat{\alpha}$], then by the continuity of the Moore-Penrose inverse [stewart1969continuity], we have
@@
$=$ $\lVert P_{\Omega^{C}}(W(1)H(1))\rVert_{F}^{2}+\lVert P_{\Omega^{C}}(\min(0,WH))\rVert_{F}^{2}-2\langle P_{\Omega^{C}}(W(1)H(1)),P_{\Omega^{C}}(\min(0,WH))\rangle$
@@
$-2\langle P_{\Omega^{C}}(\min(0,W(1)H(1))),P_{\Omega^{C}}(\min(0,WH))\rangle$
$-2\langle P_{\Omega^{C}}(\max(0,W(1)H(1))),P_{\Omega^{C}}(\min(0,WH))\rangle$
$=$ $\lVert P_{\Omega^{C}}(\min(0,W(1)H(1))-(\min(0,WH))\rVert_{F}^{2}$
$-2\langle P_{\Omega^{C}}(\max(0,W(1)H(1))),P_{\Omega^{C}}(\min(0,WH))\rangle\geq 0,$